A boat that travels 10mph in still water takes 1 hour longer to travel 24 miles against the current than it does to paddle the same distance with the current . Find the speed of the current
Answer by mananth(16946)

Boat speed = 10 mph
current speed = x mph

speed against current 10 - x
speed with current 10 + x

Distance against current 24 miles
Distance with current 24 miles

t=d/r time against = time with current -
24 / ( 10 - x )- 24 / ( 10 + x )=1

24(10+x)-24(10-x)=100-x^2
240+24x-240+24x=100-x^2
x^2+48x-100=0
x^2+50x-2x-100=0
x(x+50)-2(x+50)=0
(x+50)(x-2)=0
x= 2 OR -50
ignore negative
speed of current = 2mph